### v1

# Gentle spaghettification in effective LQG dust collapse

Abstract
Effective dust collapse inspired by loop quantum gravity predicts two main features for general realistic initial profiles: a quantum gravitational bounce of the stellar core, when the energy density becomes planckian, and shell-crossing singularities, arising within almost a planckian time after the bounce. The aim of this work is to study the mathematical and physical features of the effective spacetime near these singularities, through the Jacobi deviation equation. The results show that the radial and angular deviations between dust particles remain finite at both the bounce and shell-crossing singularity, providing a physical justification for extending the spacetime beyond shell-crossing singularities through weak solutions.
I Introduction
Spherically symmetric stellar collapse in classical general relativity generally leads to the formation of crush singularities, also called central singularities, according to the Penrose singularity theorem Penrose:1964wq . Even though such singularities have been widely studied in the literature and their mathematical properties analyzed in great detail, the common belief is that they are a pathological feature of Einstein theory of gravity, rather than a physical prediction. It has also been found that such singularities are not the only ones that develop during gravitational spherically symmetric collapse for generic initial density profiles. If neighboring layers of the collapsing star move at significantly different speeds, large inhomogeneity gradients can develop in the stellar energy density profile, sometimes leading to the so-called shell-crossing singularities (SCS) Penrose:1999vj ; Szekeres:1995gy ; Hellaby:1985zz ; Lasky:2006hq . Much less (though still significant) effort has been devoted to studying SCS compared to crush singularities, and the reasons are mainly threefold: firstly, during the collapse of pressureless matter, Hellaby and Lake proved Hellaby:1985zz that one can always choose initial data that do not develop such singularities before the central singularity is reached, which, in turn, is generally unavoidable. Additionally, by examining the mathematical features of these singularities, one can show Szekeres:1995gy that curvature scalars diverge less strongly compared to those associated with the crush singularities. Moreover, the behavior of matter falling toward these singularities is much less dramatic: despite the tidal forces being divergent at SCS, the angular and radial displacement of test particles remain finite Szekeres:1995gy . Furthermore, since the radial tidal force is negatively divergent, the so-called spaghettification effect is absent near SCS.
The purpose of this work is to understand how this picture changes once quantum gravitational corrections are taken into account. Among the possible candidate theories of quantum gravity, loop quantum gravity is one of the best-developed approaches Rovelli:2004tv ; Thiemann:2007pyv . Although a quantum gravitational model of stellar collapse is not yet available from the full theory, effective models have developed in recent years, where quantum gravitational corrections to Einstein equations are inspired by features of the underlying fundamental theory. The effective model on which this work is based is inspired from effective loop quantum cosmology, where the classical big bang singularity is replaced by a quantum gravitational bounce Ashtekar:2006wn . The modifications to Einstein field equations are based on the so-called $\bar{\mu}+K$-loop quantization scheme, where holonomy corrections with physical length $\sqrt{\Delta}$ (the minimum area gap in loop quantum gravity, proportional to the Planck length) affect only the angular component of the extrinsic curvature, and inverse triad corrections are neglected. The resulting effective dynamics has been extensively studied in recent years, both in the dust Husain:2021ojz ; Han:2023wxg ; Lewandowski:2022zce ; Giesel:2023hys ; Fazzini:2023scu ; Husain:2022gwp ; Kelly:2020lec ; Munch:2020czs ; Munch:2021oqn ; Cafaro:2024vrw ; Fazzini:2023ova ; Cipriani:2024nhx ; Bobula:2023kbo ; Bobula:2024chr ; Giesel:2023tsj and pressure cases Wilson-Ewing:2024uad ; Cafaro:2024lre . Despite much effort being devoted within this scheme to the Oppenheimer-Snyder collapse, the general picture regarding initial continuous decreasing inhomogeneous profiles is the following Fazzini:2023ova : the collapsing stellar core undergoes a quantum gravitational bounce when its energy density becomes planckian. Immediately after that, shell-crossing singularities arise, and the equations of motion break down. Although it is outside the scope of this work, for completeness, we mention that one can extend the dynamics beyond SCS formation by examining the integral form of the equation of motion, as in Painlevè-Gullstrand coordinates, it becomes a non-linear hyperbolic conservation law Leveque:2002 ; Batchelor_2000 , leading to the shockwave dynamics Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf . The same technique has been applied in the classical context Nolan:2003wp .
The question this work aims to address is how tidal forces behave during this effective dynamics, focusing in particular on the bounce point and shell-crossing singularities. The main results we found are the following: angular tidal forces remain bounded during the bounce of the stellar core and reach their maximum exactly at the bounce point. Unlike the classical case, there is a short period of time around the bounce point in which the angular tidal force is positive, and two zeroes are attained, before and after the bounce. The radial component of the tidal force, instead, is also bounded at the bouncing point, but diverges at SCS. However, unlike the classical case, where the tidal forces negatively diverges at SCS, in the effective case, the radial component of the tidal force is positive at SCS, meaning that spaghettification of a test body occurs, similarly to what happens near the classical central singularity. Despite this, importantly, by solving the radial component of the geodesic deviation equation, one finds that the radial separation between two points remains finite at the shell-crossing singularity, meaning that the spaghettification does not result in an infinite stretching of the body. This allows us to conclude that such singularities, arising during effective stellar collapse, should not be considered as dramatic as the central classical singularities, and it motivates the extension of spacetime beyond these singularities through weak solutions.
II Dust collapse in the marginally bound case
Effective dust collapse within the $\bar{\mu}+K-$loop quantization scheme has been extensively studied in Painlevè-Gullstrand coordinates Husain:2021ojz ; Kelly:2020lec ; Husain:2022gwp ; Fazzini:2023scu ; Cafaro:2024vrw ; Cipriani:2024nhx ; Fazzini:2025hsf and Lemaître-Tolman-Bondi coordinates (LTB) Giesel:2023hys ; Fazzini:2023ova ; Bobula:2024chr . The results derived in this work use LTB coordinates. The line element describing effective stellar collapse in these coordinates takes the following form
${\rm d}s^{2}=-{\rm d}t^{2}+\frac{(\partial_{R}r(R,t))^{2}}{1+\varepsilon(R)}{\rm d}R^{2}+r(R,t)^{2}{\rm d}\Omega^{2}~{},$ (1)
where $\varepsilon$ is a constant of motion fixed by the initial condition, and determines the Newtonian energy (kinetic plus potential) of any shell $R$ at infinity. We will focus here on the marginally bound case $\varepsilon(R)=0$, $\forall R$. The function $r(R,t)$ is the areal radius of each shell $R$ of the distribution at time $t$, and its explicit form is provided by the solution of the effective equations. In turn, the equation of motion in LTB coordinates takes the following form Giesel:2023hys :
$\left(\frac{\dot{r}}{r}\right)^{2}=\frac{F(R)}{r^{3}}\left(1-\frac{F(R)\gamma^{2}\Delta}{r^{3}}\right)~{},$ (2)
where we defined $F(R)\equiv 2Gm(R)$ for compactness, and $m(R)$ is the (conserved) gravitational mass within the $R-$shell. As one can explicitly notice from (2), the spacetime evolution is provided by the evolution of the areal radius of the shells composing the collapsing star. The analytic solution of the previous equation reads:
$r(R,t)=\left[F(R)\right]^{\frac{1}{3}}\left\{\frac{9}{4}\left[t-\alpha(R)\right]^{2}+\Delta\right\}^{\frac{1}{3}}~{},$ (3)
where $\alpha(R)$ is fixed by the initial condition. A fundamental feature of this generic solution, valid for any initial energy density profile, is that each shell $R$ will bounce at $t_{B}(R)=\alpha(R)$. In the pre-bounce (post-bounce) phase we have $t<\alpha(R)$ ($t>\alpha(R)$). Even though from (2) it seems that the shells’ dynamics is decoupled, and each shell evolves independently from the others (influenced only by the gravitational mass within its spherical volume), such solutions generally lead to shell-crossing singularities (SCS), at which points the equations cannot be trusted anymore. In order to define mathematically such singularities, we recall that the dust energy density in LTB coordinates reads
$\rho=\frac{\partial_{R}F}{8\pi Gr^{2}\partial_{R}r}~{}.$ (4)
A dynamical solution of the kind (3) develops SCS if, for a certain shell $R$, the following conditions are fulfilled: $\partial_{R}F\neq 0$, $\partial_{R}r(R,t)=0$. From (4), one can easily realize that these conditions imply a divergence in the dust energy density, and consequently a divergence in curvature scalars.
III Choice of initial profiles
The effective equations (2) hold for any initial energy density dust configuration in the marginally bound case. In particular, one can easily derive the effective cosmological dynamics Ashtekar:2006wn by requiring $\rho(t_{0},R)=const.$, as well as the vacuum solution Giesel:2023hys ; Giesel:2024mps , by imposing $F=const$. Here, we are interested in stellar collapse, so we will assume that the initial energy density profile is positive and of compact support. Moreover, to keep the result as general as possible, we will consider decreasing inhomogeneous profiles ($\partial_{R}\rho(t_{0},R)<0$) of compact support, without specifying the initial data.
For these kind of profiles, we can combine a theorem, stated and proved in Fazzini:2023ova , and two results, one from Fazzini:2023ova and the other from Fazzini:2025hsf :
Theorem.
For the case $\varepsilon=0$, a shell-crossing singularity forms if the initial distribution of the dust energy density $\rho(R,t_{0})$ is non-negative, continuous, of compact support, and for which m(R) is not everywhere zero,
and:
  1. 1.
    the latest time at which a shell-crossing singularity can arise in effective stellar collapse is $t=t_{B}(R)+(2/3)\sqrt{\Delta}$.
  2. 2.
    for initially decreasing continuous profiles, shell-crossing singularities cannot arise in a time $t\leq t_{B}(R)$.
Therefore, if we consider an initial profile that is non-negative, decreasing and of compact support, a shell-crossing singularity will necessarily form for a certain shell $R$ at time $t_{B}<t\leq t_{B}+(2\sqrt{\Delta}/3)$. These results allow us to restrict the analysis of tidal forces to the post-bounce dynamics of the shells, particularly to almost a planckian time interval after the bounce.
IV Geodesic deviation equation for LTB collapse
In order to understand tidal effects during generic dust collapse, we need to study the geodesic deviation equation. In the classical theory, given $u^{\mu}$ the 4-vector tangent to the geodesics, and $\delta x^{\nu}$, the displacement vector orthogonal to $u$ ($\delta x^{\nu}u_{\nu}=0$), the Jacobi equation holds Wald:1984rg
$\frac{D^{2}\delta x^{\mu}}{D\tau^{2}}=R^{\mu}_{\nu\rho\sigma}u^{\nu}u^{\rho}\delta x^{\sigma}~{},$ (5)
up to linear order in the deviation and its derivative. The covariant derivative is computed along the direction individuated by $u^{\mu}$, and therefore $\tau$ is the proper time along the geodesics. This equation does not uses the Einstein field equations but comes from the geometrical definition of the Riemann tensor. Therefore, it holds also at the effective level.
We are interested in the evolution of the deviation between dust particles during the stellar dynamics, and since dust particles follow geodesics also at the effective level (Appendix A), (5) describes tidal forces experienced by the dust itself. In particular, we are looking at a congruence of timelike geodesics for initially free-falling dust particles. Since the fluid $4$-velocity coincides with the $4$-velocity of a radially free-falling test particle, which in LTB coordinates takes the simple form $u^{\mu}=\{1,0,0,0\}$, equation (5) simplifies considerably
$\frac{D^{2}\delta x^{\mu}}{D\tau^{2}}=R^{\mu}_{00\sigma}\delta x^{\sigma}~{}.$ (6)
In order to get rid of the covariant derivative and simplify the equation further, we use a common procedure and introduce tetrads dInverno:1992gxs . The tetrads associated with the LTB line element (1), which coincide with the tetrad basis in the free-falling frame, in the marginally bound case $\varepsilon(R)=0$ take the following (classical) form:
$e^{\mu}_{0}=\left\{1,0,0,0\right\}$ (7)
$e^{\mu}_{1}=\left\{0,({\partial_{R}r})^{-1},0,0\right\}$ (8)
$e^{\mu}_{2}=\left\{0,0,r^{-1},0\right\}$ (9)
$e^{\mu}_{3}=\left\{0,0,0,({r\sin\theta})^{-1}\right\}~{}.$ (10)
These tetrads can be easily derived from $g^{\mu\nu}=e^{\mu}_{i}e^{\nu}_{j}\eta^{ij}$. Then, we write all the tensors appearing in (6) in terms of tetrads:
$R^{\mu}_{\nu\rho\sigma}=R^{a}_{bcd}e^{\mu}_{a}e^{b}_{\nu}e^{c}_{\rho}e^{d}_{\sigma}~{},$ (11)
$\delta x^{\mu}=e^{\mu}_{a}\delta x^{a}~{},$ (12)
where $a,b,c,d$ run from 00 to $3$. Notice that $\delta x^{0}=0$, since $u^{\mu}\delta x_{\mu}=0$ and $u^{\mu}$ has no spatial components. The deviation equation becomes
$\frac{D^{2}\delta x^{a}}{D\tau^{2}}=R^{a}_{00b}\delta x^{b}~{},$ (13)
which $0-$component is trivially satisfied. To arrive at (13) we used the tetrad postulate, which can be written as $De^{a}_{\mu}/D\tau=0$. The great advantage of using tetrads in LTB coordinates is that they define a locally inertial freely falling frame, for which the covariant derivative reduces to a simple ordinary derivative Szekeres:1995gy ; Fuchs(1990) . This leads to
$\frac{{\rm d}^{2}\delta x^{a}}{{\rm d}\tau^{2}}=R^{a}_{00b}\delta x^{b}~{}.$ (14)
To conclude, since the LTB coordinate time is the proper time of free-falling timelike observers, we can set $\tau=t$, and the equation further simplifies:
$\frac{\partial^{2}\delta x^{a}}{\partial t^{2}}=R^{a}_{00b}\delta x^{b}~{}.$ (15)
Although the explicit form of the effective Riemann tensor differs from the classical one, the effective LTB metric does not acquire quantum correction in its functional dependence from the solution $r(R,t)$, and the same holds for the Riemann tensor. Therefore, we can use the classical result obtained in Joshi:2024djy to write down (15) in a more explicit form:
$\partial^{2}_{t}{\delta x^{1}}=\frac{\partial^{2}_{t}\partial_{R}{r}}{\partial_{R}r}\delta x^{1}~{},$ (16)
${\partial^{2}_{t}\delta x^{i}}=\frac{\partial^{2}_{t}{r}}{r}\delta x^{i}~{},$ (17)
where the first equation determines the radial deviation and the index $i=2,3$ indicates the angular deviations in the tetradic basis. The time component of equation (15) gives a trivial identity. It is worth mentioning that, even though the functional dependence on $r$ takes the classical form, since the effective solution differs from the classical one, the same clearly holds for the solutions of equations (16) and (17).
V Tidal forces during effective stellar collapse
As pointed out before, effective inhomogeneous dust collapse does not end in the classical central singularity, which is replaced by a quantum gravitational bounce. In classical Einstein theory, angular tidal forces close to shell-crossing singularities do not diverge, while despite radial tidal forces being divergent, radial deviation does not go to zero at SCS, nor does it diverge Szekeres:1995gy . This makes such singularities less pathological than the central one. Here, we show that similar results hold also at the effective level.
To prove this, we compute explicitly (16), (17), using the solution (3) of the effective equation (2). Let’s start from the equation for angular deviation. After some computation, we find
$\partial^{2}_{t}\delta x^{i}=\frac{3\left[\Delta-\frac{3}{4}(t-\alpha)^{2}\right]}{2\left[\frac{9}{4}(t-\alpha)^{2}+\Delta\right]^{2}}\delta x^{i}~{}.$ (18)
Firstly, we notice that in the classical case $\Delta\rightarrow 0$, the angular tidal forces negatively diverges for $t\rightarrow\alpha(R)$ (classical central singularity). In contrast, here it remains always bounded, reaches its maximum at $t=\alpha(R)$ (the bouncing point), and two zeroes, for $t=\alpha\pm 2\sqrt{\Delta}/\sqrt{3}$.
The departure from the classical focusing behavior is clearly due to quantum gravity repulsion, which weakens the angular tidal force and, for almost a planckian time, produces a tidal angular repulsion.
Figure 1: Angular tidal force for $\alpha=1$, in the classical (dark blue curve) and effective collapse (red curve). In the classical case the function diverges at the crush singularity $t=\alpha=1$, while remains bounded at the effective level.
Fig.1 shows the comparison between the classical and effective cases. Notice that angular tidal forces do not diverge at shell-crossing singularity, which arises within a time $\alpha(R)<t\leq\alpha(R)+2\sqrt{\Delta}/3$.
It is worth mentioning that even though the angular tidal force remains bounded, it reaches a planckian value at the bounce, given by
$\partial^{2}_{t}\delta x^{i}|_{bounce}=\frac{3}{2\Delta}\delta x^{i}~{}.$ (19)
Let’s now examine the radial tidal force. By explicitly writing equation (16), after a lengthy but straightforward computation, we find
$\frac{\partial^{2}_{t}\partial_{R}r}{\partial_{R}r}=\left(\frac{F}{r^{3}}-\frac{10\Delta F^{2}}{r^{6}}\right)+\frac{\partial_{R}F}{\partial_{R}r}\left(\frac{4\Delta F}{r^{5}}-\frac{1}{2r^{2}}\right)~{}.$ (20)
Firstly, we notice that divergence can only occur if $\partial_{R}r=0$, provided the numerator of the last term is nonzero. This indicates that the radial tidal force remains finite at the bounce. To study the radial tidal force near and at SCS, we can assume that the first term is negligible
$\frac{\partial^{2}_{t}\partial_{R}r}{\partial_{R}r}\sim\frac{\partial_{R}F}{2r^{5}\partial_{R}r}\left(4\Delta F-r^{3}\right)~{}.$ (21)
Next, we recall that for the profiles we are considering (decreasing, continuous and of compact support), shell-crossing singularities arise for $\alpha<t\leq\alpha+2\sqrt{\Delta}/3$, which implies $F\Delta<r^{3}\leq 2F\Delta$, from (3). This means that the numerator of (21) is strictly positive, and at shell-crossing singularity the radial tidal force diverges. It is worth noting that such a divergence arises in the deep quantum gravitational regime and is a positive divergence: an infinite stretching radial force arises, conceptually similar to the one occurring classically near the central singularity, producing the spaghettification effect. In contrast, near classical shell-crossing singularities, the radial tidal force is negatively divergent: two points radially separated tend to be crushed together by the tidal force, even though in a weak fashion Szekeres:1995gy , as one can see by setting $\Delta=0$ in (21).
Let us now focus on the (positive) divergent behavior of the radial tidal force near and at the SCS. The condition for SCS can be rewritten as Fazzini:2023ova
$(t-\alpha)^{2}\partial_{R}m-2m(t-\alpha)\partial_{R}\alpha+\frac{4\Delta}{9}\partial_{R}m=0~{}.$ (22)
This is a second-degree equation in $t-\alpha$, with solutions $t_{1}=\alpha(R)+\delta t_{1}(R)$, $t_{2}=\alpha(R)+\delta t_{2}(R)$, with $\delta t_{1}\leq\delta t_{2}$. These solutions are guaranteed to be real for continuous, inhomogeneous, decreasing profiles of compact support, as shown in the previous section. It can be also easily shown that the solutions are distinct Fazzini:2023ova , except for an initial condition that satisfies $\frac{m\partial_{R}\alpha}{\partial_{R}m}=\frac{2\sqrt{\Delta}}{3}$, which is a fine-tuned case we do not consider here. Therefore, we can rewrite the condition for SCS as
$(t-\alpha(R)-\delta t_{1}(R))(t-\alpha(R)-\delta t_{2}(R))=0~{}.$ (23)
This means that close to $t_{SCS}=\alpha+\delta t_{1}$, the function $r^{\prime}$ goes to zero linearly: $r^{\prime}\sim K(t_{SCS},R)(t_{SCS}-t)$ for $t\lesssim t_{SCS}$, with $K(t_{SCS},R)>0$. Here we are only interested in tidal forces before and at the shell-crossing singularity. Beyond this point, the equation of motion in differential form breaks down, and analytic results are not reliable. With this in hand, we can rewrite (16):
$\partial^{2}_{t}\delta x^{1}\sim\frac{B(t_{SCS},R)}{t_{SCS}-t}\delta x^{1}~{},$ (24)
where $B(t_{SCS},R)$ is given by $\partial^{2}_{t}\partial_{R}r$ evaluated at shell-crossing singularity. This approximation holds only close to the SCS ($t\lesssim\alpha(R)+\delta t_{1}(R)$). Equation (24) can be solved analytically. Let us define $\delta t\equiv t_{SCS}-t>0$. The solution is
$\delta x^{1}(\delta{t})\sim$ $c_{1}\sqrt{\delta{t}}~{}I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]+$ (25)
$c_{2}\sqrt{\delta{t}}~{}K_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]~{},$ (26)
where $I_{1}$ and $K_{1}$ are modified Bessel functions of the first and second kinds, and $c_{1},c_{2}$ are constants determined by initial conditions. The previous solution holds only for $\delta{t}\rightarrow 0^{+}$. While the first term goes to zero in this limit, the second term approaches a constant non-zero value. Their explicit Taylor expansions are:
$\sqrt{\delta{t}}I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]=\sqrt{B(t_{SCS},R)}\delta{t}+O(\delta{t}^{2})~{},$ (27)
$\sqrt{\delta{t}}K_{1}\left[2\sqrt{B(R)\delta{t}}\right]=\frac{1}{2\sqrt{B(t_{SCS},R)}}+O(\delta{t})~{}.$ (28)
Therefore, even though the radial tidal forces diverge at SCS, two neighboring points that radially separated (and thus belong to two different matter layers) do not clash in general when shell-crossing singularity occur. For instance, these points can be thought of as belonging to two shells close to the ones that cross, but with a larger coordinate separation $\delta R$). If these points belong exactly to the crossing shells, they will clearly be crushed, and this condition is expected to be given by $c_{2}=0$ Szekeres:1995gy . This implies that the radial portions of the collapsing star (or even geodesic test particles radially separated) undergo spaghettification, but without being infinitely stretched. We refer to this effect as gentle spaghettification.
This result leads to the conclusion that SCS, which generally arise in effective stellar collapse, are less pathological than classical crush singularities. It also justifies the extension of the spacetime beyond SCS Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf , as done in the classical case through weak solutions. It is important to emphasize that this result is non-trivial. As mentioned before, even though the structure of the deviation equation follows the classical form, the effective solution differs from the classical one. Therefore, the behavior of tidal forces at shell-crossing singularities also differs from the classical behavior in the effective model.
To conclude, for profiles that develop shell-crossing singularities in the pre-bounce phase (where classical Einstein equations hold), the classical analysis Szekeres:1995gy can be directly exported to the effective model.
VI conclusions
Shell-crossing singularities are generally unavoidable in effective dust collapse inspired by loop quantum gravity, within the $\bar{\mu}+K$-loop quantization scheme. They therefore play a central role in such effective models, and their geometrical and physical features need to be investigated. In this work, we focused on the analysis of the tidal forces arising from effective dust collapse, with particular attention to the bouncing point and shell-crossing singularities. We found that tidal forces remain bounded at the bounce, in contrast to what happens in the classical case at the central crush singularity, and the same holds for the angular component at the shell-crossing singularity. However, the radial component of the tidal force positively diverges at the shell-crossing singularity.
Nonetheless, when studying the behavior of the radial deviation by solving the Jacobi equation, we find that it remains bounded during the whole dynamics, also when the shell-crossing singularity arises. The positive divergent behavior of the tidal force implies a spaghettification effect, similar to the one occurring near the central singularity in classical stellar collapse. However, the bounded behavior of the radial displacement between nearby points makes such an effect much less dramatic. We refer to this phenomenon as gentle spaghettification, and it allows the dynamics to extend beyond shell-crossing singularities in a meaningful way, through the integral form of the equations of motion. Future work in this direction will aim to investigate tidal forces in the effective gravitational collapse of dust in the non-marginally bound case Cipriani:2024nhx , and of fluids with pressure Cafaro:2024lre , in order to understand, at least at the numerical level, the role of pressure in the tidal forces experienced by test particles.
Acknowledgements.
This work is supported in part by the Natural Sciences and Engineering Research Council of Canada.
Appendix A Effective geodesic flow of dust particles
At the effective level the energy-momentum tensor is covariantly conserved
$\nabla^{\mu}T_{\mu\nu}=0~{}.$ (29)
This is a direct consequence of the fact that the matter part of the Einstein equation is not modified by quantum corrections. The energy-momentum tensor for dust reads $T_{\mu\nu}=\rho u_{\mu}u_{\nu}$, where $u^{\mu}$ is the fluid $4$-velocity in its rest frame. By imposing (29) one can easily get $\rho u^{\mu}\nabla_{\mu}u^{\nu}=0$, which for $\rho\neq 0$ implies that dust particles follow geodesics also at the effective level.

### v2

@@
 If neighboring layers of the collapsing star move at significantly different speeds, large inhomogeneity gradients can develop in the stellar energy density profile, sometimes leading to the so-called shell-crossing singularities (SCS) Penrose:1999vj ; Szekeres:1995gy ; Hellaby:1985zz ; Lasky:2006hq .
-Much less (though still significant) effort has been devoted to studying SCS compared to crush singularities, and the reasons are mainly threefold: firstly, during the collapse of pressureless matter, Hellaby and Lake proved Hellaby:1985zz that one can always choose initial data that do not develop such singularities before the central singularity is reached, which, in turn, is generally unavoidable.
+Although shell-crossing singularities are also physical (i.e., not mere coordinate artifacts), much less (though still significant) effort has been devoted to studying SCS compared to crush singularities, and the reasons are mainly threefold: firstly, during the collapse of pressureless matter, Hellaby and Lake proved Hellaby:1985zz that one can always choose initial data that do not develop such singularities before the central singularity is reached, which, in turn, is generally unavoidable.
 Additionally, by examining the mathematical features of these singularities, one can show Szekeres:1995gy that curvature scalars diverge less strongly compared to those associated with the crush singularities.
@@
 Although a quantum gravitational model of stellar collapse is not yet available from the full theory, effective models have developed in recent years, where quantum gravitational corrections to Einstein equations are inspired by features of the underlying fundamental theory.
-The effective model on which this work is based is inspired from effective loop quantum cosmology, where the classical big bang singularity is replaced by a quantum gravitational bounce Ashtekar:2006wn .
-The modifications to Einstein field equations are based on the so-called $\bar{\mu}+K$-loop quantization scheme, where holonomy corrections with physical length $\sqrt{\Delta}$ (the minimum area gap in loop quantum gravity, proportional to the Planck length) affect only the angular component of the extrinsic curvature, and inverse triad corrections are neglected.
-The resulting effective dynamics has been extensively studied in recent years, both in the dust Husain:2021ojz ; Han:2023wxg ; Lewandowski:2022zce ; Giesel:2023hys ; Fazzini:2023scu ; Husain:2022gwp ; Kelly:2020lec ; Munch:2020czs ; Munch:2021oqn ; Cafaro:2024vrw ; Fazzini:2023ova ; Cipriani:2024nhx ; Bobula:2023kbo ; Bobula:2024chr ; Giesel:2023tsj and pressure cases Wilson-Ewing:2024uad ; Cafaro:2024lre .
-Despite much effort being devoted within this scheme to the Oppenheimer-Snyder collapse, the general picture regarding initial continuous decreasing inhomogeneous profiles is the following Fazzini:2023ova : the collapsing stellar core undergoes a quantum gravitational bounce when its energy density becomes planckian.
-Immediately after that, shell-crossing singularities arise, and the equations of motion break down.
-Although it is outside the scope of this work, for completeness, we mention that one can extend the dynamics beyond SCS formation by examining the integral form of the equation of motion, as in Painlevè-Gullstrand coordinates, it becomes a non-linear hyperbolic conservation law Leveque:2002 ; Batchelor_2000 , leading to the shockwave dynamics Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf .
-The same technique has been applied in the classical context Nolan:2003wp .
+The effective model on which this work is based is inspired by effective loop quantum cosmology (LQC), where the classical big bang singularity is replaced by a quantum gravitational bounce Ashtekar:2006wn , arising when the energy density becomes planckian.
+Modifications to Einstein’s equations at the effective level arise from a loop quantization of the midisuperspace, with spherical symmetry imposed at the classical level.
+In LQG-inspired models, such modifications typically affect only the geometrical part of the field equations, while the matter sector remains unaltered—a strategy that is also followed in the present work.
+This approach reflects a widely adopted methodology in Loop Quantum Cosmology (LQC), where the holonomy of the extrinsic curvature is quantized—referred to as $K$-loop quantization—in contrast to the full theory, where the holonomy of the entire Ashtekar-Barbero connection is quantized Bojowald:2003mc ; Vandersloot:2006ws ; Singh_2013 .
+In this treatment, inverse triad corrections are neglected.
+The physical length over which the holonomy is computed is determined via the improved dynamics scheme (the so-called $\bar{\mu}$-scheme), which ensures that quantum corrections are suppressed in the low-curvature regime Ashtekar:2006wn and connects the effective theory to the minimum non-zero area eigenvalue of the full quantum theory $\Delta\propto l_{Pl}^{2}$.
+This quantization prescription has been implemented explicitly in Husain:2022gwp , where both the areal gauge and dust-time gauge are fixed at the classical level.
+However, the dynamics is analyzed only at the effective level.
+The expectation that these effective equations reliably capture the underlying quantum dynamics stems from an argument inherited from LQC.
+Specifically, it has been shown Rovelli:2013zaa that for cosmological solutions with physical volumes at the bounce satisfying $V_{phys}\gg\Delta^{\frac{3}{2}}$ , quantum fluctuations in the gravitational field become negligible even when the curvature approaches the Planck scale.
+This explains why quantum and effective cosmological dynamics agree so well numerically when large physical volumes are considered.
+To apply the same reasoning to gravitational collapse, consider the simplest case: Oppenheimer–Snyder (OS) collapse.
+The interior solution of the OS model is isomorphic to a cosmological spacetime, and the physical volume of the star at the bounce is much larger than Planckian Husain:2021ojz .
+Therefore, quantum fluctuations are also expected to be negligible in this astrophysical context.
+It is important to note, however, that this argument may not hold in the case of inhomogeneous collapse, where large inhomogeneities in the energy density develop—the scenario explored in this work.
+A comment on this limitation is provided in the concluding discussion.
+Motivated by feasibility and also by the previous argument, the effective dynamics has been extensively studied in recent years, both in the dust Husain:2021ojz ; Han:2023wxg ; Lewandowski:2022zce ; Giesel:2023hys ; Fazzini:2023scu ; Husain:2022gwp ; Kelly:2020lec ; Munch:2020czs ; Munch:2021oqn ; Cafaro:2024vrw ; Fazzini:2023ova ; Cipriani:2024nhx ; Bobula:2023kbo ; Bobula:2024chr ; Giesel:2023tsj and pressure cases Wilson-Ewing:2024uad ; Cafaro:2024lre .
+For a comparison with other existing effective models, see Ashtekar:2023cod .
+Despite much effort being devoted within this scheme to the Oppenheimer-Snyder collapse, where the stellar dynamics follows exactly the cosmological one Munch:2020czs ; Munch:2021oqn ; Lewandowski:2022zce ; Fazzini:2023scu ; Giesel:2023hys the general picture regarding initial continuous decreasing inhomogeneous profiles is different Fazzini:2023ova : the collapsing stellar core undergoes a quantum gravitational bounce when its energy density becomes planckian.
+Immediately after that, shell-crossing singularities arise (which, as mentioned before, are physical singularities), and the equations of motion break down.
+Shell-crossing singularities produce a non-isolated thin shell of matter, which dynamics can in principle be studied with a dynamical version of the Israel junction conditions Fazzini:2025zrq , a tool that is not yet available.
 The question this work aims to address is how tidal forces behave during this effective dynamics, focusing in particular on the bounce point and shell-crossing singularities.
-The main results we found are the following: angular tidal forces remain bounded during the bounce of the stellar core and reach their maximum exactly at the bounce point.
-Unlike the classical case, there is a short period of time around the bounce point in which the angular tidal force is positive, and two zeroes are attained, before and after the bounce.
-The radial component of the tidal force, instead, is also bounded at the bouncing point, but diverges at SCS.
-However, unlike the classical case, where the tidal forces negatively diverges at SCS, in the effective case, the radial component of the tidal force is positive at SCS, meaning that spaghettification of a test body occurs, similarly to what happens near the classical central singularity.
+The main results we found are the following: the angular tidal forces remain bounded during the bounce of the stellar core and reach their maximum exactly at the bounce point.
+Unlike the classical case, there is a short period of time around the bounce point, in which the angular tidal force is positive, and two zeroes are attained, just before and after the bounce.
+The radial component of the tidal force, instead, is also bounded at the bouncing point but diverges at SCS.
+However, unlike the classical case, where the tidal forces negatively diverge at SCS, in the effective case, the radial component of the tidal force is positive at SCS, meaning that spaghettification of a test body occurs, similarly to what happens near the classical central singularity.
 Despite this, importantly, by solving the radial component of the geodesic deviation equation, one finds that the radial separation between two points remains finite at the shell-crossing singularity, meaning that the spaghettification does not result in an infinite stretching of the body.
-This allows us to conclude that such singularities, arising during effective stellar collapse, should not be considered as dramatic as the central classical singularities, and it motivates the extension of spacetime beyond these singularities through weak solutions.
+This allows us to conclude that such singularities, arising during effective stellar collapse, should not be considered as dramatic as the central classical singularities, and it motivates the extension of spacetime beyond these physical singularities.
 II Dust collapse in the marginally bound case
 Effective dust collapse within the $\bar{\mu}+K-$loop quantization scheme has been extensively studied in Painlevè-Gullstrand coordinates Husain:2021ojz ; Kelly:2020lec ; Husain:2022gwp ; Fazzini:2023scu ; Cafaro:2024vrw ; Cipriani:2024nhx ; Fazzini:2025hsf and Lemaître-Tolman-Bondi coordinates (LTB) Giesel:2023hys ; Fazzini:2023ova ; Bobula:2024chr .
-The results derived in this work use LTB coordinates.
+Although the original formulation of the model—both at the quantum and effective levels—was carried out within the dust-time and areal gauges, it has been shown that the areal gauge condition can be relaxed at the classical level, and imposed directly at the effective level.
+This leads to an effective model featuring an unmodified classical diffeomorphism constraint, in line with the structure of the full theory Giesel:2023hys .
+This relaxation enables the study of dynamics in various spatial gauges at the effective level, with the effective equations of Husain:2022gwp recovered upon imposing the areal gauge.
+The flexibility in choosing the spatial gauge at the effective level is particularly valuable when analyzing the formation and physical properties of shell-crossing singularities Fazzini:2023ova .
+For this reason, the results presented in this work make use of LTB coordinates.
 The line element describing effective stellar collapse in these coordinates takes the following form
 ${\rm d}s^{2}=-{\rm d}t^{2}+\frac{(\partial_{R}r(R,t))^{2}}{1+\varepsilon(R)}{\rm d}R^{2}+r(R,t)^{2}{\rm d}\Omega^{2}~,$ (1)
-where $\varepsilon$ is a constant of motion fixed by the initial condition, and determines the Newtonian energy (kinetic plus potential) of any shell $R$ at infinity.
+where $\varepsilon$ is a time-independent function fixed by the initial condition, and determines the Newtonian energy (kinetic plus potential) of any shell $R$ at infinity.
 We will focus here on the marginally bound case $\varepsilon(R)=0$, $\forall R$.
-The function $r(R,t)$ is the areal radius of each shell $R$ of the distribution at time $t$, and its explicit form is provided by the solution of the effective equations.
+The function $r(R,t)$ is the areal radius of each shell $R$ of the distribution at time $t$, and its explicit form is provided by the solution of the effective equations written in the LTB gauge.
 In turn, the equation of motion in LTB coordinates takes the following form Giesel:2023hys :
 $\left(\frac{\dot{r}}{r}\right)^{2}=\frac{F(R)}{r^{3}}\left(1-\frac{F(R)\gamma^{2}\Delta}{r^{3}}\right)~,$ (2)
-where we defined $F(R)\equiv 2Gm(R)$ for compactness, and $m(R)$ is the (conserved) gravitational mass within the $R-$shell.
+where we defined $F(R)\equiv 2Gm(R)$ for compactness, and $m(R)$ is the ADM mass function for the portion of the star within the $R$-shell.
+The factor $\gamma$ is the Barbero-Immirzi parameter, that for simplicity we set to 111.
 As one can explicitly notice from (2), the spacetime evolution is provided by the evolution of the areal radius of the shells composing the collapsing star.
@@
 A fundamental feature of this generic solution, valid for any initial energy density profile, is that each shell $R$ will bounce at $t_{B}(R)=\alpha(R)$.
-In the pre-bounce (post-bounce) phase we have $t<\alpha(R)$ ($t>\alpha(R)$).
-Even though from (2) it seems that the shells’ dynamics is decoupled, and each shell evolves independently from the others (influenced only by the gravitational mass within its spherical volume), such solutions generally lead to shell-crossing singularities (SCS), at which points the equations cannot be trusted anymore.
-In order to define mathematically such singularities, we recall that the dust energy density in LTB coordinates reads
-$\rho=\frac{\partial_{R}F}{8\pi Gr^{2}\partial_{R}r}~{}.$ (4)
-A dynamical solution of the kind (3) develops SCS if, for a certain shell $R$, the following conditions are fulfilled: $\partial_{R}F\neq 0$, $\partial_{R}r(R,t)=0$.
-From (4), one can easily realize that these conditions imply a divergence in the dust energy density, and consequently a divergence in curvature scalars.
+In the pre-bounce (post-bounce) phase, we have $t<\alpha(R)$ ($t>\alpha(R)$).
+Even though from (2) it seems that the shells’ dynamics is decoupled, and each shell evolves independently from the others (influenced only by the gravitational mass within its spherical volume), such solutions generally lead to shell-crossing singularities, at which point the equations cannot be trusted anymore.
+In order to define mathematically such singularities, we recall that the dust energy density in LTB coordinates reads Giesel:2023hys
+$\rho=\frac{\mathcal{H}_{dust}}{4\pi\sqrt{q}}=\frac{\partial_{R}F}{8\pi Gr^{2}\partial_{R}r}~.$ (4)
+Here, $\sqrt{q}$ denotes the determinant of the spatial metric (1), and $\mathcal{H}_{dust}$ is the dust Hamiltonian density, integrated over ${\rm d}\theta$ and ${\rm d}\phi$, while $\rho$ represents the dust energy density.
+Note that these expressions retain their classical functional form, since the matter sector is kept classical in this effective model Kelly:2020lec .
+As in the classical case, the LTB coordinates are adapted to the comoving dust frame.
+Therefore, $\rho$ corresponds to the $tt$ component of the energy-momentum tensor.
+Note that the functional relation between the lagrangian density for dust and the energy-momentum tensor is classical.
+Although the functional form of (4) remains classical, its explicit behavior deviates from the classical solution when $\rho$ approaches the critical density $\rho_{crit.}=3/(8\pi G\Delta)\sim\rho_{\text{Pl.}}$, or equivalently when $r(R,t)\sim[F(R)\Delta]^{1/3}$—that is, near the bounce.
+A dynamical solution of the form (3) develops a SCS if, for a given shell labeled by $R$, the following conditions are satisfied: $\partial_{R}F\neq 0$ and $\partial_{R}r(R,t)=0$.
+From equation (4), it is easy to see that these conditions lead to a divergence in the dust energy density, which in turn results in a divergence of the curvature scalars Giesel:2024mps .
 III Choice of initial profiles
@@
 In particular, one can easily derive the effective cosmological dynamics Ashtekar:2006wn by requiring $\rho(t_{0},R)=const.$, as well as the vacuum solution Giesel:2023hys ; Giesel:2024mps , by imposing $F=const$.
-Here, we are interested in stellar collapse, so we will assume that the initial energy density profile is positive and of compact support.
+Here, we are interested in stellar collapse, so we will assume that the initial energy density profile is positive and has compact support.
 Moreover, to keep the result as general as possible, we will consider decreasing inhomogeneous profiles ($\partial_{R}\rho(t_{0},R)<0$) of compact support, without specifying the initial data.
-For these kind of profiles, we can combine a theorem, stated and proved in Fazzini:2023ova , and two results, one from Fazzini:2023ova and the other from Fazzini:2025hsf :
+For these kinds of profile, we can combine a theorem, stated and proved in Fazzini:2023ova , and two results, one from Fazzini:2023ova and the other from Fazzini:2025hsf :
 Theorem.
@@
   1. 1.
-    the latest time at which a shell-crossing singularity can arise in effective stellar collapse is $t=t_{B}(R)+(2/3)\sqrt{\Delta}$.
+    the latest time at which a shell-crossing singularity can arise in effective stellar collapse is $t=t_{B}(R)+(2/3)\sqrt{\Delta}$ Fazzini:2023ova .
   2. 2.
-    for initially decreasing continuous profiles, shell-crossing singularities cannot arise in a time $t\leq t_{B}(R)$.
-Therefore, if we consider an initial profile that is non-negative, decreasing and of compact support, a shell-crossing singularity will necessarily form for a certain shell $R$ at time $t_{B}<t\leq t_{B}+(2\sqrt{\Delta}/3)$.
+    for initially decreasing continuous profiles, shell-crossing singularities cannot arise in a time $t\leq t_{B}(R)$ Fazzini:2025hsf .
+Therefore, if we consider an initial profile that is non-negative, decreasing, and of compact support, a shell-crossing singularity will necessarily form for a certain shell $R$ at time $t_{B}<t\leq t_{B}+(2\sqrt{\Delta}/3)$.
 These results allow us to restrict the analysis of tidal forces to the post-bounce dynamics of the shells, particularly to almost a planckian time interval after the bounce.
 IV Geodesic deviation equation for LTB collapse
-In order to understand tidal effects during generic dust collapse, we need to study the geodesic deviation equation.
+In order to understand the tidal effects during generic dust collapse, we need to study the geodesic deviation equation.
 In the classical theory, given $u^{\mu}$ the 4-vector tangent to the geodesics, and $\delta x^{\nu}$, the displacement vector orthogonal to $u$ ($\delta x^{\nu}u_{\nu}=0$), the Jacobi equation holds Wald:1984rg
@@
 The covariant derivative is computed along the direction individuated by $u^{\mu}$, and therefore $\tau$ is the proper time along the geodesics.
-This equation does not uses the Einstein field equations but comes from the geometrical definition of the Riemann tensor.
-Therefore, it holds also at the effective level.
-We are interested in the evolution of the deviation between dust particles during the stellar dynamics, and since dust particles follow geodesics also at the effective level (Appendix A), (5) describes tidal forces experienced by the dust itself.
-In particular, we are looking at a congruence of timelike geodesics for initially free-falling dust particles.
-Since the fluid $4$-velocity coincides with the $4$-velocity of a radially free-falling test particle, which in LTB coordinates takes the simple form $u^{\mu}=\{1,0,0,0\}$, equation (5) simplifies considerably
+This equation does not use the field equations (classical or modified), but comes from the geometrical definition of the Riemann tensor.
+Therefore, it also holds at the effective level.
+We are interested in the evolution of the deviation between dust particles during the stellar dynamics, and since dust particles follow geodesics also at the effective level (which can be readily verified by checking that the dust four-velocity $u^{\mu}=(1,0,0,0)$ satisfies the geodesic equation), (5) describes tidal forces experienced by the dust itself.
+In particular, we are looking at a congruence of time-like geodesics for initially free-falling dust particles.
+Since the fluid 444-velocity coincides with the 444-velocity of a radially free-falling test particle, equation (5) simplifies considerably
 $\frac{D^{2}\delta x^{\mu}}{D\tau^{2}}=R^{\mu}_{00\sigma}\delta x^{\sigma}~.$ (6)
-In order to get rid of the covariant derivative and simplify the equation further, we use a common procedure and introduce tetrads dInverno:1992gxs .
+In order to get rid of the covariant derivative and further simplify the equation, we use a common procedure and introduce tetrads dInverno:1992gxs .
 The tetrads associated with the LTB line element (1), which coincide with the tetrad basis in the free-falling frame, in the marginally bound case $\varepsilon(R)=0$ take the following (classical) form:
@@
 $e^{\mu}_{3}=\left\{0,0,0,({r\sin\theta})^{-1}\right\}~.$ (10)
-These tetrads can be easily derived from $g^{\mu\nu}=e^{\mu}_{i}e^{\nu}_{j}\eta^{ij}$. Then, we write all the tensors appearing in (6) in terms of tetrads:
+These tetrads can easily be derived from $g^{\mu\nu}=e^{\mu}_{i}e^{\nu}_{j}\eta^{ij}$. Then, we write all the tensors appearing in (6) in terms of tetrads:
 $R^{\mu}_{\nu\rho\sigma}=R^{a}_{bcd}e^{\mu}_{a}e^{b}_{\nu}e^{c}_{\rho}e^{d}_{\sigma}~,$ (11)
 $\delta x^{\mu}=e^{\mu}_{a}\delta x^{a}~,$ (12)
-where $a,b,c,d$ run from 00 to $3$. Notice that $\delta x^{0}=0$, since $u^{\mu}\delta x_{\mu}=0$ and $u^{\mu}$ has no spatial components. The deviation equation becomes
+where $a,b,c,d$ run from 0 to 333. Notice that $\delta x^{0}=0$, since $u^{\mu}\delta x_{\mu}=0$ and $u^{\mu}$ have no spatial components. The deviation equation becomes
 $\frac{D^{2}\delta x^{a}}{D\tau^{2}}=R^{a}_{00b}\delta x^{b}~,$ (13)
-which $0-$component is trivially satisfied.
+which 00-0 -component is trivially satisfied.
 To arrive at (13) we used the tetrad postulate, which can be written as $De^{a}_{\mu}/D\tau=0$.
@@
 where the first equation determines the radial deviation and the index $i=2,3$ indicates the angular deviations in the tetradic basis.
-The time component of equation (15) gives a trivial identity.
-It is worth mentioning that, even though the functional dependence on $r$ takes the classical form, since the effective solution differs from the classical one, the same clearly holds for the solutions of equations (16) and (17).
+The time component of (15) gives a trivial identity.
+It is worth mentioning that even though the functional dependence on $r$ takes the classical form, since the effective solution differs from the classical one, the same clearly holds for the solutions of equations (16) and (17).
 V Tidal forces during effective stellar collapse
 As pointed out before, effective inhomogeneous dust collapse does not end in the classical central singularity, which is replaced by a quantum gravitational bounce.
-In classical Einstein theory, angular tidal forces close to shell-crossing singularities do not diverge, while despite radial tidal forces being divergent, radial deviation does not go to zero at SCS, nor does it diverge Szekeres:1995gy .
+In classical Einstein theory, the angular tidal forces close to shell-crossing singularities do not diverge, while despite radial tidal forces being divergent, the radial deviation does not go to zero at SCS, nor does it diverge Szekeres:1995gy .
 This makes such singularities less pathological than the central one.
-Here, we show that similar results hold also at the effective level.
-To prove this, we compute explicitly (16), (17), using the solution (3) of the effective equation (2). Let’s start from the equation for angular deviation. After some computation, we find
+Here, we show that similar results also hold at the effective level.
+To prove this, we compute explicitly (16), (17), using the solution (3) of the effective equation (2). Let us start from the equation for the angular deviation. After some computation, we find
 $\partial^{2}_{t}\delta x^{i}=\frac{3\left[\Delta-\frac{3}{4}(t-\alpha)^{2}\right]}{2\left[\frac{9}{4}(t-\alpha)^{2}+\Delta\right]^{2}}\delta x^{i}~.$ (18)
-Firstly, we notice that in the classical case $\Delta\rightarrow 0$, the angular tidal forces negatively diverges for $t\rightarrow\alpha(R)$ (classical central singularity).
-In contrast, here it remains always bounded, reaches its maximum at $t=\alpha(R)$ (the bouncing point), and two zeroes, for $t=\alpha\pm 2\sqrt{\Delta}/\sqrt{3}$.
+Firstly, we note that in the classical case $\Delta\rightarrow 0$, the angular tidal forces negatively diverge for $t\rightarrow\alpha(R)$ (classical central singularity).
+In contrast, here it remains always bounded, reaches its maximum at $t=\alpha(R)$ (the bouncing point) and has two zeros, for $t=\alpha\pm 2\sqrt{\Delta}/\sqrt{3}$.
 The departure from the classical focusing behavior is clearly due to quantum gravity repulsion, which weakens the angular tidal force and, for almost a planckian time, produces a tidal angular repulsion.
 Figure 1: Angular tidal force for $\alpha=1$, in the classical (dark blue curve) and effective collapse (red curve).
+The dotted vertical line signals the latest time at which a SCS occurs in the effective case.
 In the classical case the function diverges at the crush singularity $t=\alpha=1$, while remains bounded at the effective level.
 Fig.1 shows the comparison between the classical and effective cases.
-Notice that angular tidal forces do not diverge at shell-crossing singularity, which arises within a time $\alpha(R)<t\leq\alpha(R)+2\sqrt{\Delta}/3$.
+Notice that angular tidal forces do not diverge at shell-crossing singularity.
+Although the latest time at which a SCS arises and the largest zero of (18) are quantitatively similar, they must be considered as two distinct features.
 It is worth mentioning that even though the angular tidal force remains bounded, it reaches a planckian value at the bounce, given by
 $\partial^{2}_{t}\delta x^{i}|_{bounce}=\frac{3}{2\Delta}\delta x^{i}~.$ (19)
-Let’s now examine the radial tidal force. By explicitly writing equation (16), after a lengthy but straightforward computation, we find
+Let us now examine the radial tidal force.
+By explicitly writing the right-hand side of equation (16) (divided by $\delta x^{1}$), after a lengthy but straightforward computation, we find (see Appendix A)
 $\frac{\partial^{2}_{t}\partial_{R}r}{\partial_{R}r}=\left(\frac{F}{r^{3}}-\frac{10\Delta F^{2}}{r^{6}}\right)+\frac{\partial_{R}F}{\partial_{R}r}\left(\frac{4\Delta F}{r^{5}}-\frac{1}{2r^{2}}\right)~.$ (20)
-Firstly, we notice that divergence can only occur if $\partial_{R}r=0$, provided the numerator of the last term is nonzero.
-This indicates that the radial tidal force remains finite at the bounce.
+Firstly, we notice that divergence can only occur if $\partial_{R}r=0$, provided that the numerator of the last term is non-zero.
+This indicates that the radial tidal force remains finite at the bounce, while it instead diverges at SCS.
 To study the radial tidal force near and at SCS, we can assume that the first term is negligible
 $\frac{\partial^{2}_{t}\partial_{R}r}{\partial_{R}r}\sim\frac{\partial_{R}F}{2r^{5}\partial_{R}r}\left(4\Delta F-r^{3}\right)~.$ (21)
-Next, we recall that for the profiles we are considering (decreasing, continuous and of compact support), shell-crossing singularities arise for $\alpha<t\leq\alpha+2\sqrt{\Delta}/3$, which implies $F\Delta<r^{3}\leq 2F\Delta$, from (3).
-This means that the numerator of (21) is strictly positive, and at shell-crossing singularity the radial tidal force diverges.
+Next, we recall that for the profiles we are considering (decreasing, continuous, and of compact support), shell-crossing singularities arise for $\alpha<t\leq\alpha+2\sqrt{\Delta}/3$, which implies $F\Delta<r^{3}\leq 2F\Delta$, from (3).
+This means that the numerator of (21) is strictly positive and that at the shell-crossing singularity the radial tidal force diverges.
 It is worth noting that such a divergence arises in the deep quantum gravitational regime and is a positive divergence: an infinite stretching radial force arises, conceptually similar to the one occurring classically near the central singularity, producing the spaghettification effect.
 In contrast, near classical shell-crossing singularities, the radial tidal force is negatively divergent: two points radially separated tend to be crushed together by the tidal force, even though in a weak fashion Szekeres:1995gy , as one can see by setting $\Delta=0$ in (21).
-Let us now focus on the (positive) divergent behavior of the radial tidal force near and at the SCS. The condition for SCS can be rewritten as Fazzini:2023ova
-$(t-\alpha)^{2}\partial_{R}m-2m(t-\alpha)\partial_{R}\alpha+\frac{4\Delta}{9}\partial_{R}m=0~{}.$ (22)
+Let us now focus on the (positive) divergent behavior of the radial tidal force near and at the SCS. The condition for SCS can be derived from the solution (3) and rewritten as Fazzini:2023ova
+$(t-\alpha)^{2}\partial_{R}m-2m(t-\alpha)\partial_{R}\alpha+\frac{4\Delta}{9}\partial_{R}m=0~,$ (22)
+see also Appendix B.
 This is a second-degree equation in $t-\alpha$, with solutions $t_{1}=\alpha(R)+\delta t_{1}(R)$, $t_{2}=\alpha(R)+\delta t_{2}(R)$, with $\delta t_{1}\leq\delta t_{2}$.
-These solutions are guaranteed to be real for continuous, inhomogeneous, decreasing profiles of compact support, as shown in the previous section.
-It can be also easily shown that the solutions are distinct Fazzini:2023ova , except for an initial condition that satisfies $\frac{m\partial_{R}\alpha}{\partial_{R}m}=\frac{2\sqrt{\Delta}}{3}$, which is a fine-tuned case we do not consider here.
+These solutions are guaranteed to be real and positive for continuous, inhomogeneous, decreasing profiles (for which $\partial_{R}\alpha>0$ Fazzini:2025hsf ) of compact support, as a consequence of the results in the previous section.
+It can also be easily shown that the solutions are distinct Fazzini:2023ova , except for an initial condition that satisfies $\frac{m\partial_{R}\alpha}{\partial_{R}m}=\frac{2\sqrt{\Delta}}{3}$, which is a fine-tuned case which we do not consider here.
 Therefore, we can rewrite the condition for SCS as
 $(t-\alpha(R)-\delta t_{1}(R))(t-\alpha(R)-\delta t_{2}(R))=0~.$ (23)
-This means that close to $t_{SCS}=\alpha+\delta t_{1}$, the function $r^{\prime}$ goes to zero linearly: $r^{\prime}\sim K(t_{SCS},R)(t_{SCS}-t)$ for $t\lesssim t_{SCS}$, with $K(t_{SCS},R)>0$.
-Here we are only interested in tidal forces before and at the shell-crossing singularity.
-Beyond this point, the equation of motion in differential form breaks down, and analytic results are not reliable.
+This means that close to $t=t_{1}=\alpha+\delta t_{1}$, the function $r^{\prime}$ goes to zero linearly: $\partial_{R}r\sim K(t_{1},t_{2},R)(t_{1}-t)$ for $t\lesssim t_{1}$, with $K(t_{1},t_{2},R)>0$ (see Appendix B).
+Here we are only interested in the tidal forces before and at the shell-crossing singularity.
+Beyond this point, the analytic solution (3) is not reliable.
 With this in hand, we can rewrite (16):
-$\partial^{2}_{t}\delta x^{1}\sim\frac{B(t_{SCS},R)}{t_{SCS}-t}\delta x^{1}~{},$ (24)
-where $B(t_{SCS},R)$ is given by $\partial^{2}_{t}\partial_{R}r$ evaluated at shell-crossing singularity.
-This approximation holds only close to the SCS ($t\lesssim\alpha(R)+\delta t_{1}(R)$).
+$\partial^{2}_{t}\delta x^{1}\sim\frac{B(t_{1},t_{2},R)}{t_{1}-t}\delta x^{1}~,$ (24)
+where $B(t_{1},t_{2},R)$ is given by $\partial^{2}_{t}\partial_{R}r$ evaluated at shell-crossing singularity divided by $K(t_{1},t_{2},R)$.
+It is important to keep in mind that this approximation holds only close to the SCS ($t\lesssim\alpha(R)+\delta t_{1}(R)$).
 Equation (24) can be solved analytically.
-Let us define $\delta t\equiv t_{SCS}-t>0$.
+Let us call $t_{1}=t_{SCS}$, and define $\delta t\equiv t_{SCS}-t>0$.
 The solution is
-$\delta x^{1}(\delta{t})\sim$ $c_{1}\sqrt{\delta{t}}~{}I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]+$ (25)
-$c_{2}\sqrt{\delta{t}}~{}K_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]~{},$ (26)
-where $I_{1}$ and $K_{1}$ are modified Bessel functions of the first and second kinds, and $c_{1},c_{2}$ are constants determined by initial conditions.
+$\delta x^{1}(\delta{t})\sim$ $c_{1}(R)\sqrt{\delta{t}}~I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]+$ (25)
+$c_{2}(R)\sqrt{\delta{t}}~K_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]~,$ (26)
+where $I_{1}$ and $K_{1}$ are modified Bessel functions of the first and second kinds, and $c_{1}(R),c_{2}(R)$ are time independent functions determined by initial conditions.
 The previous solution holds only for $\delta{t}\rightarrow 0^{+}$.
@@
 $\sqrt{\delta{t}}K_{1}\left[2\sqrt{B(R)\delta{t}}\right]=\frac{1}{2\sqrt{B(t_{SCS},R)}}+O(\delta{t})~.$ (28)
-Therefore, even though the radial tidal forces diverge at SCS, two neighboring points that radially separated (and thus belong to two different matter layers) do not clash in general when shell-crossing singularity occur.
-For instance, these points can be thought of as belonging to two shells close to the ones that cross, but with a larger coordinate separation $\delta R$).
-If these points belong exactly to the crossing shells, they will clearly be crushed, and this condition is expected to be given by $c_{2}=0$ Szekeres:1995gy .
+Therefore, even though the radial tidal forces diverge at SCS, two neighboring points that are radially separated (and thus belong to two different matter layers) do not generally clash when a shell-crossing singularity occurs.
+For example, these points can be thought of as belonging to two shells close to the ones that cross, but with a larger coordinate separation $\delta R$).
+If these points belong exactly to the crossing shells, they will be crushed, and this condition is given by $c_{2}(R)=0$ Szekeres:1995gy , since at SCS $\partial_{R}r(R,t_{SCS})=0$, that implies $\delta r=\delta x^{1}=0$.
 This implies that the radial portions of the collapsing star (or even geodesic test particles radially separated) undergo spaghettification, but without being infinitely stretched.
 We refer to this effect as gentle spaghettification.
-This result leads to the conclusion that SCS, which generally arise in effective stellar collapse, are less pathological than classical crush singularities.
-It also justifies the extension of the spacetime beyond SCS Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf , as done in the classical case through weak solutions.
-It is important to emphasize that this result is non-trivial.
-As mentioned before, even though the structure of the deviation equation follows the classical form, the effective solution differs from the classical one.
+This result leads to the conclusion that SCSs, which generally arise in effective stellar collapse, are less pathological than classical crush singularities.
+It also justifies the extension of the spacetime beyond SCS Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf ; Liu:2025fil ; Fazzini:2025zrq ; Husain:2025wrh .
+It is important to emphasize that this result is nontrivial.
+As mentioned above, even though the structure of the deviation equation follows the classical form, the effective solution differs from the classical one.
 Therefore, the behavior of tidal forces at shell-crossing singularities also differs from the classical behavior in the effective model.
-To conclude, for profiles that develop shell-crossing singularities in the pre-bounce phase (where classical Einstein equations hold), the classical analysis Szekeres:1995gy can be directly exported to the effective model.
+To conclude, for profiles that develop shell-crossing singularities in the pre-bounce phase (where the classical Einstein equations hold), the classical analysis Szekeres:1995gy can be directly exported to the effective model.
 VI conclusions
-Shell-crossing singularities are generally unavoidable in effective dust collapse inspired by loop quantum gravity, within the $\bar{\mu}+K$-loop quantization scheme.
-They therefore play a central role in such effective models, and their geometrical and physical features need to be investigated.
-In this work, we focused on the analysis of the tidal forces arising from effective dust collapse, with particular attention to the bouncing point and shell-crossing singularities.
+Shell-crossing singularities are physical singularities generally unavoidable in effective dust collapse inspired by loop quantum gravity, within the $\bar{\mu}+K$-loop quantization scheme.
+They therefore play a central role in this effective model, and their geometrical and physical features need to be investigated.
+In this work, we focus on the analysis of the tidal forces arising from effective dust collapse, with particular attention to the bouncing point and shell-crossing singularities.
 We found that tidal forces remain bounded at the bounce, in contrast to what happens in the classical case at the central crush singularity, and the same holds for the angular component at the shell-crossing singularity.
 However, the radial component of the tidal force positively diverges at the shell-crossing singularity.
-Nonetheless, when studying the behavior of the radial deviation by solving the Jacobi equation, we find that it remains bounded during the whole dynamics, also when the shell-crossing singularity arises.
+Nevertheless, when studying the behavior of the radial deviation by solving the Jacobi equation, we find that it remains bounded throughout the dynamics, also when the shell-crossing singularity arises.
 The positive divergent behavior of the tidal force implies a spaghettification effect, similar to the one occurring near the central singularity in classical stellar collapse.
@@
 We refer to this phenomenon as gentle spaghettification, and it allows the dynamics to extend beyond shell-crossing singularities in a meaningful way, through the integral form of the equations of motion.
-Future work in this direction will aim to investigate tidal forces in the effective gravitational collapse of dust in the non-marginally bound case Cipriani:2024nhx , and of fluids with pressure Cafaro:2024lre , in order to understand, at least at the numerical level, the role of pressure in the tidal forces experienced by test particles.
+Future work in this direction will aim to investigate the tidal forces in the effective gravitational collapse of dust in the non-marginally bound case Cipriani:2024nhx , and of fluids with pressure Cafaro:2024lre , in order to understand, at least at the numerical level, the role of pressure in the tidal forces experienced by test particles.
+Although this work does not aim to study the spacetime evolution beyond the formation of shell-crossing singularities, the analysis presented here could be applied to spacetimes where such an extension has been performed, particularly to investigate tidal forces near the non-isolated thin shell generated by the SCS.
+When SCSs develop, the effective equations expressed in Painlevé-Gullstrand (PG) coordinates exhibit characteristic crossing, which imply a multi-valued behavior of the extrinsic curvature.
+Since these equations take the form of hyperbolic conservation laws in the marginally bound case Husain:2022gwp and balance laws in the non-marginally bound case Cipriani:2024nhx , one might be tempted to study their dynamics through weak solutions, as commonly done in other fields where characteristic crossings occur Batchelor ; Hoefer2006 .
+This approach results in an evolving shock in the extrinsic curvature Husain:2022gwp ; Cipriani:2024nhx .
+However, it was recently shown Fazzini:2025zrq that the integral formulation underlying weak solutions implicitly assumes the continuity of PG time throughout the shock (the SCS in LTB coordinates), which is physically questionable.
+Indeed, the Misner-Sharp mass function becomes discontinuous at the shock, implying different time dilations measured by interior and exterior metrics, and consequently different PG times flowing.
+Furthermore, weak solutions exhibit superluminal behavior.
+The proper way to extend the spacetime beyond the SCS is to apply the Israel junction conditions to a non-isolated thin shell, ensuring subluminal propagation and allowing PG time to be discontinuous at the shock.
+Once this framework is established, understanding the evolution of tidal forces near the evolving thin shell will be crucial.
+To conclude, a natural question arises as to whether shell-crossing singularities can be avoided once a full quantum theory of dust collapse is taken into account, going beyond the effective approach.
+In principle, quantum gravitational fluctuations in the fields could introduce corrections to the effective equation (2) that prevent the formation of SCS and lead to a singularity-free evolution, although this outcome is not guaranteed.
+If this scenario holds, the thin shell dynamics beyond SCS studied here might qualitatively capture the behavior predicted by a more refined effective model.
+Moreover, since quantum fluctuations are expected to decay rapidly near the SCS, the tidal forces computed in this work could provide a reasonable approximation to those in such a hypothetical effective model.
 Acknowledgements.
 This work is supported in part by the Natural Sciences and Engineering Research Council of Canada.
-Appendix A Effective geodesic flow of dust particles
-At the effective level the energy-momentum tensor is covariantly conserved
-$\nabla^{\mu}T_{\mu\nu}=0~{}.$ (29)
-This is a direct consequence of the fact that the matter part of the Einstein equation is not modified by quantum corrections.
-The energy-momentum tensor for dust reads $T_{\mu\nu}=\rho u_{\mu}u_{\nu}$, where $u^{\mu}$ is the fluid $4$-velocity in its rest frame.
-By imposing (29) one can easily get $\rho u^{\mu}\nabla_{\mu}u^{\nu}=0$, which for $\rho\neq 0$ implies that dust particles follow geodesics also at the effective level.
+Appendix A The radial tidal force
+We can compute the quantity $\frac{\partial_{t}^{2}\partial_{R}r}{\partial_{R}r}$ from the general solution (3).
+By deriving twice this expression in time, and rewriting the result in term of $r(R,t)$, we get
+$\partial_{t}^{2}r(R,t)=\frac{F(R)^{\frac{2}{3}}}{r^{4}}\left[r^{3}+2\Delta F(R)\right]~.$ (29)
+Then, deriving with respect of the radial coordinate, one gets
+$\partial_{R}{\partial_{t}^{2}r(R,t)}=\frac{2\Delta}{r^{4}}\partial_{R}F^{\frac{5}{3}}+\frac{\partial_{R}F^{\frac{2}{3}}}{r}-F^{\frac{2}{3}}\left(1+\frac{8\Delta F}{r^{3}}\right)\frac{\partial_{R}r}{r^{2}}~.$ (30)
+By dividing this expression by $\partial_{R}r(R,t)$, and rearranging terms, we get
+$\frac{\partial^{2}_{t}\partial_{R}r}{\partial_{R}r}=\left(\frac{F}{r^{3}}-\frac{10\Delta F^{2}}{r^{6}}\right)+\frac{\partial_{R}F}{\partial_{R}r}\left(\frac{4\Delta F}{r^{5}}-\frac{1}{2r^{2}}\right)~.$ (31)
+Appendix B Behavior of $\partial_{R}r(R,t)$ close to SCS
+In this appendix, we prove that, close to SCS $t\lesssim t_{1}$, we can write $\partial_{R}r(R,t)=K(t_{1},t_{2},R)(t_{1}-t)+O(t_{1}-t)$, where $t_{1}(R)$ and $t_{2}(R)$ are the two real roots of (22), and $K(t_{1},t_{2},R)$ is a positive function.
+We can start by differentiating (3) with respect to $R$, getting
+$\frac{\partial_{R}r}{r}=\frac{\partial_{R}m}{3m}+\frac{3}{2}\frac{(\alpha(R)-t)\partial_{R}\alpha}{\left[\frac{9}{4}(\alpha-t)^{2}+\Delta\right]}~,$ (32)
+which can be rewritten as
+$\partial_{r}(R,t)=\frac{3r}{4m}\frac{\partial_{R}m(\alpha-t)^{2}+\frac{4}{9}\Delta\partial_{R}m-2\partial_{R}\alpha(t-\alpha)}{\frac{9}{4}(t-\alpha)^{2}+\Delta}~.$ (33)
+We recognize the last term in the numerator as the left-hand side of (22), and therefore we can rewrite the previous equation as
+$\partial_{R}r(R,t)=\frac{3r}{4m}\frac{[t-t_{1}(R)][t-t_{2}(R)]}{\frac{9}{4}(t-\alpha)^{2}+\Delta}~,$ (34)
+where $t_{1}(R)$ and $t_{2}(R)$ are the solutions of (22).
+Now, let us study this expression for $t-t_{1}=-\varepsilon$, $\varepsilon>0$ (close but before SCS).
+By explicitly writing $r(R,t)$ using (3), we get
+$\partial_{R}r(R,t_{1}-\varepsilon)=\frac{3F^{\frac{1}{3}}}{4m(R)}\frac{\varepsilon(\varepsilon+t_{2}(R)-t_{1}(R))}{\left[\frac{9}{4}(t_{1}-\alpha(R)-\varepsilon)^{2}+\Delta\right]^{\frac{2}{3}}}~,$ (35)
+and, for $\varepsilon<<1$:
+$\partial_{R}r(R,t_{1}-\varepsilon)=\frac{3F^{\frac{1}{3}}}{4m(R)}\frac{\varepsilon(t_{2}(R)-t_{1}(R))}{\left[\frac{9}{4}(t_{1}-\alpha(R))^{2}+\Delta\right]^{\frac{2}{3}}}+O(\varepsilon)~.$ (36)
+If we define
+$K(t_{1},t_{2},R)\equiv\frac{3F}{4m}\frac{(t_{2}(R)-t_{1}(R))}{\left[\frac{9}{4}(t_{1}-\alpha(R))^{2}+\Delta\right]^{\frac{2}{3}}}$ (37)
+then we can write
+$\partial_{R}r(R,t)=K(t_{1},t_{2},R)(t_{1}-t)+O(t_{1}-t)~,$ (38)
+for $t\lesssim t_{1}$, with $K(t_{1},t_{2},R)>0$.